# Diffusion State-Guided Projected Gradient for Inverse Problems

### v1

@@ C.4 ReSample
Our DiffStateGrad-ReSample algorithm integrates the state-guided projected gradient into the optimization process of ReSample (Song et al., 2023a). We introduce two new hyperparameters: the variance retention threshold $\tau$ and a frequency $F$ for applying our DiffStateGrad step. During each ReSample step, we first perform SVD on the current latent representation (or diffusion state) ${\bm{Z}}^{\prime}_{t}$ (${\bm{z}}^{\prime}_{t}$ in image matrix form). Note that we do not perform SVD within the gradient descent loop itself, meaning that we only perform SVD at most once per iteration of the sampling algorithm. We then determine the appropriate rank based on $\tau$ and construct projection matrices. Then, within the gradient descent loop for solving $\hat{{\bm{z}}}_{0}({\bm{y}})$, we approximate the gradient in the diffusion state subspace using our projection matrices every $F=5$ steps. On steps where DiffStateGrad is not applied, we use the standard gradient. This adaptive, periodic application of DiffStateGrad allows for a balance between the benefits of low-rank approximation and the potential need for full gradient information. The rest of the ReSample algorithm, including the stochastic resampling step, remains unchanged.
@@
1:Measurements ${\bm{y}}$, $\mathcal{A}(\cdot)$, Encoder $\mathcal{E}(\cdot)$, Decoder $\mathcal{D}(\cdot)$, Score function ${\bm{s}}_{\theta}(\cdot,t)$, Pretrained LDM Parameters $\beta_{t}$, $\bar{\alpha}_{t}$, $\eta$, $\delta$, Hyperparameter $\gamma$ to control $\sigma_{t}^{2}$, Time steps to perform resample $C$, Variance retention threshold $\tau$, Frequency $F$
@@
17:              if  step number $\mod F=0$ then
@@
C.5 DAPS
We improve upon DAPS by incorporating a state-guided projected gradient. We introduce a variance retention threshold $\tau$ to determine the projection rank. For each noise level in the annealing loop, DAPS computes the initial estimate $\hat{{\bm{x}}}_{0}^{(0)}$ by solving the probability flow ODE using the score model ${\bm{s}}_{\theta}$. This estimate represents a guess of the clean image given the current noisy sample ${\bm{x}}_{t_{i}}$. We then perform SVD on this estimate in image matrix form $\hat{\bm{X}}_{0}^{(0)}$ (using it as our diffusion state), determine the appropriate rank based on $\tau$, and construct projection matrices. Within the Langevin dynamics loop, we calculate the full gradient, combining both the prior term $\log p(\hat{{\bm{x}}}_{0}^{(j)}|{\bm{x}}_{t_{i}})$ and the likelihood term $\log p({\bm{y}}|\hat{{\bm{x}}}_{0}^{(j)})$. For each step (i.e., frequency $F=1$), we project this gradient using our pre-computed matrices and use this projected gradient for the update step. This process is repeated for each noise level, progressively refining our estimate of the clean image.

### v2

@@ C.4 ReSample
Our DiffStateGrad-ReSample algorithm integrates the state-guided projected gradient into the optimization process of ReSample (Song et al., 2023a). We introduce two new hyperparameters: the variance retention threshold $\tau$ and a period $P$ for applying our DiffStateGrad step. During each ReSample step, we first perform SVD on the current latent representation (or diffusion state) ${\bm{Z}}^{\prime}_{t}$ (${\bm{z}}^{\prime}_{t}$ in image matrix form). Note that we do not perform SVD within the gradient descent loop itself, meaning that we only perform SVD at most once per iteration of the sampling algorithm. We then determine the appropriate rank based on $\tau$ and construct projection matrices. Then, within the gradient descent loop for solving $\hat{{\bm{z}}}_{0}({\bm{y}})$, we approximate the gradient in the diffusion state subspace using our projection matrices every $P=5$ steps. On steps where DiffStateGrad is not applied, we use the standard gradient. This adaptive, periodic application of DiffStateGrad allows for a balance between the benefits of low-rank approximation and the potential need for full gradient information. The rest of the ReSample algorithm, including the stochastic resampling step, remains unchanged.
@@
1:Measurements ${\bm{y}}$, $\mathcal{A}(\cdot)$, Encoder $\mathcal{E}(\cdot)$, Decoder $\mathcal{D}(\cdot)$, Score function ${\bm{s}}_{\theta}(\cdot,t)$, Pretrained LDM Parameters $\beta_{t}$, $\bar{\alpha}_{t}$, $\eta$, $\delta$, Hyperparameter $\gamma$ to control $\sigma_{t}^{2}$, Time steps to perform resample $C$, Variance retention threshold $\tau$, Period $P$
@@
17:              if  step number $\mod P=0$ then
@@
C.5 DPS
Like DiffStateGrad-PSLD, our DiffStateGrad-DPS algorithm integrates the state-guided projected gradient into each iteration of the DPS process. The key difference lies in operating directly on pixel-based diffusion states rather than latent representations. After computing the standard DPS update ${\bm{x}}^{\prime}_{t-1}$, we follow the same projection strategy: performing SVD on the current diffusion state ${\bm{X}}_{t}$, determining projection rank using threshold $\tau$, and constructing an approximated gradient $\bm{G}^{\prime}_{t}$ that replaces the standard DPS gradient update.
For experiments, we use the official implementation of DPS (Chung et al., 2023) with default configurations.
Algorithm 4 DiffStateGrad-DPS for Image Restoration Tasks
1:$T,\bm{y},\{\zeta_{t}\}_{t=1}^{T},\{\tilde{\sigma}_{t}\}_{t=1}^{T},\bm{s}_{\theta},$variance retention threshold $\tau$
2:${\bm{x}}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$
3:for $t=T-1$ to 00 do
4:     $\hat{\bm{s}}\leftarrow\bm{s}_{\theta}({\bm{x}}_{t},t)$
5:     $\hat{{\bm{x}}}_{0}\leftarrow\frac{1}{\sqrt{\bar{\alpha_{t}}}}({\bm{x}}_{t}+(1-\bar{\alpha_{t}})\hat{\bm{s}})$
6:     ${\bm{z}}\sim\mathcal{N}(\bm{0},\bm{I})$
7:     ${\bm{x}}^{\prime}_{t-1}\leftarrow\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}{\bm{x}}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\hat{{\bm{x}}}_{0}+\tilde{\sigma}_{t}{\bm{z}}$
8:     $\bm{g}_{t}\leftarrow\zeta_{t}\nabla_{{\bm{x}}_{t}}\|\bm{y}-\mathcal{A}(\hat{{\bm{x}}}_{0})\|^{2}_{2}$
9:     $\bm{U},\bm{S},\bm{V}\leftarrow\text{SVD}(\bm{X}_{t})$
10:     $\lambda_{j}\leftarrow s_{j}^{2}$ ($s_{j}$ are the singular values of $\bm{S}$)
11:     $c_{k}\leftarrow\frac{\sum_{j=1}^{k}\lambda_{j}}{\sum_{j}\lambda_{j}}$
[…]
13:     $\bm{A}_{t}\leftarrow\bm{U}_{r}$
14:     $\bm{B}_{t}\leftarrow\bm{V}_{r}$
15:     $\bm{R}_{t}\leftarrow\bm{A}_{t}^{T}\bm{G}_{t}\bm{B}_{t}^{T}$
16:     $\bm{G}^{\prime}_{t}\leftarrow\bm{A}_{t}\bm{R}_{t}\bm{B}_{t}$
17:     ${\bm{x}}_{t-1}\leftarrow{\bm{x}}^{\prime}_{t-1}-\bm{g}^{\prime}_{t}$
18:end for
19:return $\hat{{\bm{x}}}_{0}$
C.6 DAPS
We improve upon DAPS by incorporating a state-guided projected gradient. We introduce a variance retention threshold $\tau$ to determine the projection rank. For each noise level in the annealing loop, DAPS computes the initial estimate $\hat{{\bm{x}}}_{0}^{(0)}$ by solving the probability flow ODE using the score model ${\bm{s}}_{\theta}$. This estimate represents a guess of the clean image given the current noisy sample ${\bm{x}}_{t_{i}}$. We then perform SVD on this estimate in image matrix form $\hat{\bm{X}}_{0}^{(0)}$ (using it as our diffusion state), determine the appropriate rank based on $\tau$, and construct projection matrices. Within the Langevin dynamics loop, we calculate the full gradient, combining both the prior term $\log p(\hat{{\bm{x}}}_{0}^{(j)}|{\bm{x}}_{t_{i}})$ and the likelihood term $\log p({\bm{y}}|\hat{{\bm{x}}}_{0}^{(j)})$. For each step (i.e., period $P=1$), we project this gradient using our pre-computed matrices and use this projected gradient for the update step. This process is repeated for each noise level, progressively refining our estimate of the clean image.
@@ C.6 DAPS
C.7 DDNM
We report numbers from DAPS (Zhang et al., 2024), which conducts experiments under identical settings to ours.
C.8 DDRM
We report numbers from DAPS (Zhang et al., 2024), which conducts experiments under identical settings to ours.
C.9 MCG
We report numbers from DPS (Chung et al., 2023), which conducts experiments under identical settings to ours.
C.10 MPGD-AE
For experiments, we use the official implementation of MPGD-AE (He et al., 2024). For accurate comparison, we use 1000 DDIM steps and the guidance weight parameter of 0.5 (which we find through fine-tuning).
@@ Appendix D Proofs
Appendix E MRI Experiments
To demonstrate the applicability of DiffStateGrad beyond natural images and the discussed diffusion-based inverse problems, we conduct an additional experiment on Magnetic Resonance Imaging (MRI); this represents an inverse problem where undersampled measurements are observed in the frequency domain, and the task is to reconstruct high-quality MRI in the image space. We utilize the Compressed Sensing Generative Model  (Jalal et al., 2021), which employs Langevin dynamics for MRI reconstruction (available at https://github.com/utcsilab/csgm-mri-langevin). This method, which we refer to as CSGM-Langevin, is built upon prior works (Song & Ermon, 2020; 2019).
We incorporate our proposed DiffStateGrad into the measurement gradient guidance. Specifically, given the complex nature of MRI data, we apply DiffStateGrad to the magnitude of the complex-valued data while preserving the phase information. This framework uses ${\bm{x}}_{t}$ for both measurement gradient computation and incorporation. The unconditional diffusion model was trained on T2-weighted brain datasets from the NYU fastMRI dataset (Zbontar et al., 2018; Knoll et al., 2020), and the reported results were averaged over $30$ test examples (reporting avg (std)). The measurement operators give samples vertically equispaced in $k$-space at an undersampling rate of $R=4$.
While the default optimized measurement gradient step size from (Jalal et al., 2021) is $5.0$, the results confirm that DiffStateGrad enhances the robustness of CSGM-Langevin when the step size deviates from this optimal value. For example, CSGM-Langevin’s performance is significantly reduced when the measurement gradient step size increases to $7.5$ or $10$, while DiffStateGrad’s performance decline is less. This suggests that DiffStateGrad can be a practical tool to reduce the risk of drastic failure in real-world applications. Finally, given the small sample size and the preliminary nature of this analysis, we leave extensive analysis for future work.
(a) PSNR.
(b) SSIM.
(c) LPIPS.
Figure 19: Robustness of DiffStateGrad to MG step size in MRI. Comparing CSGM-Langevin (Jalal et al., 2021) with DiffStateGrad-CSGM-Langevin on $30$ examples from the test set of the T2-weighted brain fastMRI dataset (Zbontar et al., 2018; Knoll et al., 2020). The measurement operator gives vertically and equispaced samples in the $k$-space with the undersampling rate of $R=4$.
E.1 NYU fastMRI dataset
The data used in the above experiment was obtained from the NYU fastMRI Initiative database (fastmri.med.nyu.edu(Zbontar et al., 2018; Knoll et al., 2020). We note that NYU fastMRI investigators provided data but did not participate in the analysis or writing of this report. See their website for more information on the primary goal of fastMRI to advance machine learning research for the reconstruction of medical images.